OpenStudy (blackops2luvr):

A cylindrical rain barrel has a radius of 2 feet and holds a total of 30 cubic feet of water. How tall is the rain barrel? Use 3.14 for pi. Round your answer to the nearest hundredth. 1.58 feet 2.39 feet 3.57 feet 4.78 feet @AnswerMyQuestions @iGreen

OpenStudy (igreen):

\(\pi r^2 h\) 30 = (3.14)(2^2)h 30 = (3.14)(4)h 30 = 12.56h 30 / 12.56 = ?

OpenStudy (anonymous):

I think you multiply 3.14 x 2^2 first, then divide 30 by that.

OpenStudy (igreen):

@blackops2luvr

OpenStudy (anonymous):

Yup, 30 / 12.56

OpenStudy (blackops2luvr):

2.38853503185

OpenStudy (anonymous):

Yup.

OpenStudy (blackops2luvr):

B ?

OpenStudy (igreen):

Yep.